# Efficient Spin-Adapted Implementation of Multireference Algebraic Diagrammatic Construction Theory. I. Core-Ionized States and X-Ray Photoelectron Spectra

### v1

@@ Abstract
We present an efficient implementation of multireference algebraic diagrammatic construction theory (MR-ADC) for simulating core-ionized states and X-ray photoelectron spectra (XPS). Taking advantage of spin adaptation, automatic code generation, and density fitting, our implementation can perform calculations for molecules with more than 1500 molecular orbitals, incorporating static and dynamic correlation in the ground and excited electronic states. We demonstrate the capabilities of MR-ADC methods by simulating the XPS spectra of substituted ferrocene complexes and azobenzene isomers. For the ground electronic states of these molecules, the XPS spectra computed using the extended second-order MR-ADC method (MR-ADC(2)-X) are in excellent agreement with available experimental results. We further show that MR-ADC can be used as a tool for interpreting or predicting the results of time-resolved XPS measurements by simulating the core ionization spectra of azobenzene along its photoisomerization, including the XPS signatures of excited states and the minimum energy conical intersection. This work is the first in a series of publications reporting the efficient implementations of MR-ADC methods.
@@ 1 Introduction
Here, we demonstrate the capabilities of our efficient MR-ADC implementation by simulating and interpreting the XPS spectra of substituted ferrocene complexes in their ground electronic states and the azobenzene molecule along its excited-state isomerization. Our calculations employ large core-valence polarized basis sets correlating all electrons in up to 1532 molecular orbitals. For the substituted ferrocene complexes, we show that the carbon K-edge XPS spectra simulated using the extended second-order MR-ADC method (MR-ADC(2)-X) are in excellent agreement with the experimental data, providing accurate interpretation of overlapping spectral features. In our study of azobenzene, we predict the carbon and nitrogen K-edge XPS signatures of ground and excited electronic states at the cis, trans, and conical intersection geometries, which may be helpful in interpreting the results of TR-XPS experiments in the future.
@@ 2.2 MR-ADC for core-ionized states and X-ray photoelectron spectra
To overcome this challenge, IP-MR-ADC is combined with a core-valence separation (CVS)70, 71 approximation that neglects the coupling between valence- and core-ionized states due to their large difference in energy and spatial localization.39, 40 In the CVS-IP-MR-ADC methods, the molecular orbitals of the system are split into four subsets: “core”, “valence”, “active”, and “virtual” (Figure 1). The union of core and valence subspaces corresponds to the doubly occupied non-active orbitals in the reference CASSCF wavefunction . The core subspace includes all lowest-energy molecular orbitals starting with the one that is expected to be ionized first in the XPS spectrum. The remaining non-active occupied orbitals are incorporated in the valence subspace. The CVS-IP-MR-ADC matrices (, , and ) are expressed in the basis of ionized electronic configurations where the th-order excitation operators are required to ionize or excite an electron from at least one core orbital (Figure 1).
@@ 2.3 Spin-free formulation of CVS-IP-MR-ADC
When using a non-relativistic Hamiltonian, it is possible and computationally advantageous to remove the dependence on spin in equations through so-called spin adaptation. One of the most widely employed approaches to spin-adapt multireference theories72, 73 is to express all operators in terms of spin-free unitary group generators (e.g., ) that are invariant under the SU(2) transformations of spin-orbitals where and indicate the up and down spin, respectively. Once the operators are defined, their matrix elements are derived using the spin-free formulation of Wick’s theorem, yielding fully spin-adapted equations.74 While this approach can be straightforwardly used in state-specific multireference methods, it is less convenient for multistate effective Hamiltonian theories such as MR-ADC where the excited-state spin symmetry or particle number can be different from that of the reference electronic state.
In this work, we employ the spin adaptation approach developed by Kutzelnigg, Shamasundar, and Mukherjee75, 76, 72, 73 (KSM) that allows to eliminate spin variables from spin-orbital equations a posteriori, by utilizing the relationships between spin-free and spin-dependent tensors for the reference state. Importantly, the resulting equations can be used to perform spin-adapted calculations starting with a closed- () or open-shell () reference state, as long as the reference wavefunction is an equally weighted ensemble of the entire spin multiplet (, ). Such reference state can be computed using the state-averaged CASSCF method, which is available in many quantum chemistry software packages. In addition to its straightforward implementation, the KSM approach enforces the degeneracy of computed open-shell excited states, which can be violated in calculations with a pure-state open-shell reference. The KSM spin adaptation was successfully used to develop efficient implementations of several effective Hamiltonian theories, such as state-specific partially internally contracted multireference coupled cluster theory,77, 78 multireference equation-of-motion coupled cluster theory,79, 80 anti-Hermitian contracted Schrödinger equation,81, 82, 83, 84 and multireference driven similarity renormalization group.85
@@
Additionally, , , and depend on up to four-particle active-space reduced density matrices of (-RDM, ). The relationships between spin-orbital () and spin-free () RDMs up to have been previously reported.73 For the 1- and 2-RDMs, these equations can be written as:
For , these relationships are rather complicated and can be found in the Supplementary Information.73 To the best of our knowledge, the equations for spin-adapting 4-RDM have not been published. We derived them using the approach outlined in Ref. 73 and included in the Supplementary Information. We note that the IP-MR-ADC(2) and IP-MR-ADC(2)-X methods can be implemented without 4-RDM by factorizing its contributions into intermediates,38 although we have not taken advantage of this yet in our efficient implementation.
@@ 3.1 Overview
The CVS-IP-MR-ADC(2) and CVS-IP-MR-ADC(2)-X methods were implemented using Python in the open-source and freely available program Prism.86 To obtain the CASSCF orbitals, one- and two-electron integrals, and the reference reduced density matrices, Prism was interfaced with the PySCF package.87 The CVS-IP-MR-ADC code follows an algorithm summarized below:
@@
Our efficient implementation of CVS-IP-MR-ADC(2) and CVS-IP-MR-ADC(2)-X has several features: (i) spin adaption of all tensor contractions (Section 2.3) assisted by automatic equation and code generation as discussed in Section 3.2; (ii) efficient in-core and out-of-core handling of two-electron integrals utilizing density fitting (Section 3.3) and h5py library;89 (iii) optimized implementation of tensor contractions using basic linear algebra subroutines (BLAS), opteinsum90 and numpy91 modules; and (iv) open multiprocessing (OpenMP) parallelization of computationally intensive tasks. In the following, we provide additional details on automatic equation and code generation (Section 3.2) and the use of density fitting (Section 3.3).
@@ 3.2 Automatic Derivation of Spin-Free Equations and Code Generation
Computer-aided derivation and code generation enabled many significant advances in quantum chemistry. Widespread adoption of second quantization formalism92, 93 created a need for tools that can handle tedious operations, such as permuting strings of creation and annihilation operators, normal ordering, evaluating commutators, and matrix elements. Automatic derivation of second-quantized equations was pioneered by Janssen and Schaefer in the framework of coupled cluster theory (CC).94 Since then, implementations of many single-95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113 and multireference methods114, 79, 78, 115, 116, 117, 118 have been developed with the help of computer tools for equation and code generation, such as tensor contraction engine (TCE),97 SecondQuantizationAlgebra (SQA),119 SMITH,104 SMITH3,120 ORCA-AGE,121 sympy,122 ,123 WICK&D,124 and QCMATH.125, 126
In our spin-orbital implementation of MR-ADC methods, we utilized the SQA program119 to derive expressions for the elements of , , , and their matrix-vector products.37, 35, 38 In this work, we developed an extension of SQA called SQA+127 that allows to perform spin adaptation using the KSM approach discussed in Section 2.3. In SQA+, every operator is initially represented in a spin-orbital form with separate labels for molecular orbitals and electron spin. The spatial orbitals are classified according to one of the four subspaces shown in Figure 1. Once an operator expression is defined by the user, its matrix elements are evaluated in the spin-orbital form by normal-ordering and () with respect to 37 and discarding vanishing terms. The spin-orbital tensor contractions are converted to the spin-free form using the KSM relationships from Section 2.3 and Supplementary Information, and are simplified by combining identical terms. The resulting expressions are converted into Python code, which can be directly incorporated into Prism. Additionally, SQA+ allows to rewrite the spin-free equations in terms of automatically defined reusable intermediates,35 which can be used to make the code more efficient. We note that an implementation of spin-free unitary group generators in SQA has been developed by Saitow et al.,115 although we did not employ it in our work.
@@ 3.3 Density Fitting
To enable the CVS-IP-MR-ADC calculations with large one-electron basis sets, our efficient implementation utilized the density fitting (DF) approximation128, 129, 130, 131, 132
@@
The DF approximation lowers the cost of transforming two-electron integrals from atomic to molecular basis and allows to avoid storing on disk, which becomes prohibitively expensive in calculations with more than 1000 molecular orbitals. Density fitting has been widely used to reduce the cost of excited-state electronic structure methods.133, 134, 135, 136, 137, 138, 139, 140, 141, 142 Here, we employ it for the first time in the implementation of MR-ADC.
To assess the accuracy of DF approximation in the CVS-IP-MR-ADC calculations, we computed core ionization energies for molecules from the benchmark set of Liu et al.143. The results presented in the Supplementary Information demonstrate that using the exact in the reference CASSCF calculations and DF-approximated two-electron integrals in CVS-IP-MR-ADC has a very small effect on core ionization energies with mean absolute error (MAE) of eV. Introducing the DF approximation in the CASSCF step increases the MAE to 0.0011 eV. Overall, these results suggest that DF can be used to significantly lower the cost of CVS-IP-MR-ADC calculations without compromising their accuracy.
@@ 4 Computational Details
The ground-state geometries of all molecules were optimized using density functional theory (DFT)144, 145, 146, 147 implemented in the Molpro148, 149, 150 package. To compute the – MECI geometry of azobenzene, we employed the time-dependent DFT (TD-DFT)151, 152, 153 algorithm154, 155 implemented in Q-Chem.156 All DFT and TD-DFT calculations used the B3LYP hybrid exchange-correlation functional,157, 158, 159, 160 the D3(BJ) dispersion correction,161 and the cc-pVQZ basis set with JK density fitting.162, 163, 164
@@
The core ionization energies and XPS spectra were computed using the efficient implementation of MR-ADC in the Prism program.86 As discussed in Section 3, Prism was interfaced with PySCF to obtain the CASSCF orbitals, one- and two-electron integrals, and reference reduced density matrices. All MR-ADC calculations employed the correlation-consistent core-polarized cc-pwCVQZ basis set for carbon and nitrogen168, 169, 170 and the cc-pVQZ basis for hydrogen and iron. Density fitting was used for the reference CASSCF and excited-state MR-ADC methods utilizing the cc-pwCVQZ-JKFIT162, 164 and cc-pwCVQZ-RIFIT162, 171 auxiliary basis sets, respectively. The MR-ADC calculations were performed using the and parameters to remove linearly dependent semiinternal and double excitations, respectively.37 Scalar relativistic effects were incorporated using the spin-free exact-two-component (X2C) Hamiltonian.172, 173, 143, 174, 175 The experimental spectra were digitized using WebPlotDigitizer.176
@@ 5.1 Substituted Ferrocenes
We first apply the efficient MR-ADC implementation developed in this work to three ferrocene derivatives with the ethyl (Et), vinyl (V), or ethynyl (E) substituent (R) in one of the cyclopentadienyl (Cp) rings (Figure 2). The redox potentials of these molecules are highly sensitive to the R group, making them attractive surface modifiers for metals and semiconductors in memory devices, electrochemical sensors, batteries, and fuel cells.178, 179, 180, 181, 182 Meanwhile, the properties of ferrocene and its derivatives are known to be strongly influenced by electron correlation effects and have been the focus of many theoretical studies.183, 184, 185, 186, 187, 188, 189, 190, 191, 192, 193, 194, 195, 196 The XPS spectra of EtFC, VFC, and EFC in the gas phase have been measured experimentally,177 allowing us to test the performance of MR-ADC for these challenging systems.
Figure 2a shows the experimental carbon K-edge XPS spectrum of ethyl-ferrocene (EtFC)177 along with the theoretical spectra calculated at the MR-ADC(2) and MR-ADC(2)-X levels of theory. The experimental XPS spectrum shows a broad feature with a maximum at 290.0 eV and a shoulder around 290.5 eV. The MR-ADC(2) and MR-ADC(2)-X calculations reproduce the experimental spectrum quite well when supplied with 0.20 eV broadening. The best agreement with the experiment is shown by MR-ADC(2)-X, which underestimates the band maximum by 0.4 eV and accurately describes the relative energy of the shoulder feature. MR-ADC(2) overestimates the band maximum by 1.4 eV and exhibits a somewhat more pronounced shoulder.
@@
and is consistent with the enhanced -delocalization of CASSCF natural orbitals (Figure S3). MR-ADC(2) overestimates the peak position and spacing relative to the experiment, predicting a weak shoulder feature with the selected broadening parameter. This discrepancy can be attributed to the overestimation of core ionization energy for the terminal carbon atom of the vinyl group and is likely associated with the less accurate description of excited-state orbital relaxation effects in MR-ADC(2) compared to MR-ADC(2)-X.
Finally, we consider the experimental177 and simulated C K-edge XPS spectra of ethynyl-ferrocene (EFC) shown in Figure 2c. Similar to EtFC, the EFC core ionization spectra exhibit a strong peak followed by a shoulder feature 0.5 eV higher in energy. As for EtFC and VFC, MR-ADC(2)-X has the best agreement with the experiment, slightly underestimating the shoulder intensity. In contrast to EtFC where the shoulder was assigned to the C atom of substituent, the MR-ADC(2)-X calculations suggest that the shoulder in the EFC spectrum originates from the Cp carbon directly bonded to the ethynyl group. These results suggest that ethynyl acts as an acceptor, depleting the electron density on the Cp carbon adjacent to the substituent and weakening the screening of corresponding core-ionized state. Indeed, the analysis of CASSCF natural orbitals (Figure S4) shows a significant occupation of in-plane and out-of-plane -antibonding ethynyl orbitals, indicating noticeable back-donation from Cp to the ethynyl group. MR-ADC(2) shows a larger error in the peak maximum and predicts a more intense shoulder than what is observed in the experiment as a result of overestimating the relative energy of core-hole states in the substituent.
Overall, our results demonstrate that the efficient MR-ADC implementation developed in this work can routinely simulate the K-edge XPS spectra of transition metal complexes with large basis sets. Specifically, the MR-ADC(2) and MR-ADC(2)-X calculations reported here were performed correlating all electrons in 1412 (EFC), 1472 (VFC), and 1532 (EtFC) molecular orbitals. Using a single Intel Xeon Gold 6148 computer node with 40 CPUs, the wall time of MR-ADC(2) and MR-ADC(2)-X simulations did not exceed 4 and 40 hours, respectively, after completing the reference CASSCF calculations. Consistent with our earlier benchmarks,39, 40 the MR-ADC(2)-X results are in excellent agreement with the experimental gas-phase XPS spectra.

### v2

@@ Abstract
We present an efficient implementation of multireference algebraic diagrammatic construction theory (MR-ADC) for simulating core-ionized states and X-ray photoelectron spectra (XPS). Taking advantage of spin adaptation, automatic code generation, and density fitting, our implementation can perform calculations for molecules with more than 1500 molecular orbitals, incorporating static and dynamic correlation in the ground and excited electronic states. We demonstrate the capabilities of MR-ADC methods by simulating the XPS spectra of substituted ferrocene complexes and azobenzene isomers. For the ground electronic states of these molecules, the XPS spectra computed using the extended second-order MR-ADC method (MR-ADC(2)-X) are in a very good agreement with available experimental results. We further show that MR-ADC can be used as a tool for interpreting or predicting the results of time-resolved XPS measurements by simulating the core ionization spectra of azobenzene along its photoisomerization, including the XPS signatures of excited states and the minimum energy conical intersection. This work is the first in a series of publications reporting the efficient implementations of MR-ADC methods.
@@ 1 Introduction
Here, we demonstrate the capabilities of our efficient MR-ADC implementation by simulating and interpreting the XPS spectra of substituted ferrocene complexes in their ground electronic states and the azobenzene molecule along its excited-state isomerization. Our calculations employ large core-valence polarized basis sets correlating all electrons in up to 1532 molecular orbitals. For the substituted ferrocene complexes, we show that the carbon K-edge XPS spectra simulated using the extended second-order MR-ADC method (MR-ADC(2)-X) are in a very good agreement with the experimental data, providing accurate interpretation of overlapping spectral features. In our study of azobenzene, we predict the carbon and nitrogen K-edge XPS signatures of ground and excited electronic states at the cis, trans, and conical intersection geometries, which may be helpful in interpreting the results of TR-XPS experiments in the future.
@@ 2.2 MR-ADC for core-ionized states and X-ray photoelectron spectra
To overcome this challenge, IP-MR-ADC is combined with a core-valence separation (CVS)70, 71, 72, 73, 74 approximation that neglects the coupling between valence- and core-ionized states due to their large difference in energy and spatial localization.39, 40 In the CVS-IP-MR-ADC methods, the molecular orbitals of the system are split into four subsets: “core”, “valence”, “active”, and “virtual” (Figure 1). The union of core and valence subspaces corresponds to the doubly occupied non-active orbitals in the reference CASSCF wavefunction . The core subspace includes all lowest-energy molecular orbitals starting with the one that is expected to be ionized first in the XPS spectrum. The remaining non-active occupied orbitals are incorporated in the valence subspace. The CVS-IP-MR-ADC matrices (, , and ) are expressed in the basis of ionized electronic configurations where the th-order excitation operators are required to ionize or excite an electron from at least one core orbital (Figure 1).
@@
We note that the CVS approximation used in this work is different from the one originally proposed in the context of single-reference ADC methods72, 73, 74 where double excitations from two core orbitals were excluded. Including these doubly excited configurations has been shown to significantly improve the accuracy of CVS approximation75, 76, 77 and has been widely used in other electronic structure methods based on CVS78, 79, 80, 81, 82, 83, 84.
@@ 2.3 Spin-free formulation of CVS-IP-MR-ADC
When using a non-relativistic Hamiltonian, it is possible and computationally advantageous to remove the dependence on spin in equations through so-called spin adaptation. One of the most widely employed approaches to spin-adapt multireference theories85, 86 is to express all operators in terms of spin-free unitary group generators (e.g., ) that are invariant under the SU(2) transformations of spin-orbitals where and indicate the up and down spin, respectively. Once the operators are defined, their matrix elements are derived using the spin-free formulation of Wick’s theorem, yielding fully spin-adapted equations.87 While this approach can be straightforwardly used in state-specific multireference methods, it is less convenient for multistate effective Hamiltonian theories such as MR-ADC where the excited-state spin symmetry or particle number can be different from that of the reference electronic state.
In this work, we employ the spin adaptation approach developed by Kutzelnigg, Shamasundar, and Mukherjee88, 89, 85, 86 (KSM) that allows to eliminate spin variables from spin-orbital equations a posteriori, by utilizing the relationships between spin-free and spin-dependent tensors for the reference state. Importantly, the resulting equations can be used to perform spin-adapted calculations starting with a closed- () or open-shell () reference state, as long as the reference wavefunction is an equally weighted ensemble of the entire spin multiplet (, ). Such reference state can be computed using the state-averaged CASSCF method, which is available in many quantum chemistry software packages. In addition to its straightforward implementation, the KSM approach enforces the degeneracy of computed open-shell excited states, which can be violated in calculations with a pure-state open-shell reference. The KSM spin adaptation was successfully used to develop efficient implementations of several effective Hamiltonian theories, such as state-specific partially internally contracted multireference coupled cluster theory,90, 91 multireference equation-of-motion coupled cluster theory,92, 93 anti-Hermitian contracted Schrödinger equation,94, 95, 96, 97 and multireference driven similarity renormalization group.98
@@
Additionally, , , and depend on up to four-particle active-space reduced density matrices of (-RDM, ). The relationships between spin-orbital () and spin-free () RDMs up to have been previously reported.86 For the 1- and 2-RDMs, these equations can be written as:
For , these relationships are rather complicated and can be found in the Supplementary Information.86 To the best of our knowledge, the equations for spin-adapting 4-RDM have not been published. We derived them using the approach outlined in Ref. 86 and included in the Supplementary Information. We note that the IP-MR-ADC(2) and IP-MR-ADC(2)-X methods can be implemented without 4-RDM by factorizing its contributions into intermediates,38 although we have not taken advantage of this yet in our efficient implementation.
@@ 3.1 Overview
The CVS-IP-MR-ADC(2) and CVS-IP-MR-ADC(2)-X methods were implemented using Python in the open-source and freely available program Prism.99 To obtain the CASSCF orbitals, one- and two-electron integrals, and the reference reduced density matrices, Prism was interfaced with the PySCF package.100 The CVS-IP-MR-ADC code follows an algorithm summarized below:
@@
Our efficient implementation of CVS-IP-MR-ADC(2) and CVS-IP-MR-ADC(2)-X has several features: (i) spin adaption of all tensor contractions (Section 2.3) assisted by automatic equation and code generation as discussed in Section 3.2; (ii) efficient in-core and out-of-core handling of two-electron integrals utilizing density fitting (Section 3.3) and h5py library;102 (iii) optimized implementation of tensor contractions using basic linear algebra subroutines (BLAS), opteinsum103 and numpy104 modules; and (iv) open multiprocessing (OpenMP) parallelization of computationally intensive tasks. In the following, we provide additional details on automatic equation and code generation (Section 3.2) and the use of density fitting (Section 3.3).
@@ 3.2 Automatic Derivation of Spin-Free Equations and Code Generation
Computer-aided derivation and code generation enabled many significant advances in quantum chemistry. Widespread adoption of second quantization formalism105, 106 created a need for tools that can handle tedious operations, such as permuting strings of creation and annihilation operators, normal ordering, evaluating commutators, and matrix elements. Automatic derivation of second-quantized equations was pioneered by Janssen and Schaefer in the framework of coupled cluster theory (CC).107 Since then, implementations of many single-108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125 and multireference methods126, 127, 92, 91, 128, 129, 130, 131 have been developed with the help of computer tools for equation and code generation, such as tensor contraction engine (TCE),110 SecondQuantizationAlgebra (SQA),132 SMITH,117 SMITH3,133 ORCA-AGE,134 sympy,135 ,136 WICK&D,137 and QCMATH.138, 139
In our spin-orbital implementation of MR-ADC methods, we utilized the SQA program132 to derive expressions for the elements of , , , and their matrix-vector products.37, 35, 38 In this work, we developed an extension of SQA called SQA+140 that allows to perform spin adaptation using the KSM approach discussed in Section 2.3. In SQA+, every operator is initially represented in a spin-orbital form with separate labels for molecular orbitals and electron spin. The spatial orbitals are classified according to one of the four subspaces shown in Figure 1. Once an operator expression is defined by the user, its matrix elements are evaluated in the spin-orbital form by normal-ordering and () with respect to 37 and discarding vanishing terms. The spin-orbital tensor contractions are converted to the spin-free form using the KSM relationships from Section 2.3 and Supplementary Information, and are simplified by combining identical terms. The resulting expressions are converted into Python code, which can be directly incorporated into Prism. Additionally, SQA+ allows to rewrite the spin-free equations in terms of automatically defined reusable intermediates,35 which can be used to make the code more efficient. We note that an implementation of spin-free unitary group generators in SQA has been developed by Saitow et al.,128 although we did not employ it in our work.
@@ 3.3 Density Fitting
To enable the CVS-IP-MR-ADC calculations with large one-electron basis sets, our efficient implementation utilized the density fitting (DF) approximation141, 142, 143, 144, 145
@@
The DF approximation lowers the cost of transforming two-electron integrals from atomic to molecular basis and allows to avoid storing on disk, which becomes prohibitively expensive in calculations with more than 1000 molecular orbitals. Density fitting has been widely used to reduce the cost of excited-state electronic structure methods.146, 147, 148, 149, 150, 151, 152, 153, 154, 155 Here, we employ it for the first time in the implementation of MR-ADC.
To assess the accuracy of DF approximation in the CVS-IP-MR-ADC calculations, we computed core ionization energies for molecules from the benchmark set of Liu et al.156. The results presented in the Supplementary Information demonstrate that using the exact in the reference CASSCF calculations and DF-approximated two-electron integrals in CVS-IP-MR-ADC has a very small effect on core ionization energies with mean absolute error (MAE) of eV. Introducing the DF approximation in the CASSCF step increases the MAE to 0.0011 eV. Overall, these results suggest that DF can be used to significantly lower the cost of CVS-IP-MR-ADC calculations without compromising their accuracy.
@@ 4 Computational Details
The ground-state geometries of all molecules were optimized using density functional theory (DFT)157, 158, 159, 160 with the B3LYP hybrid exchange-correlation functional161, 162, 163, 164 and the D3(BJ) dispersion correction.165 To compute the – MECI geometry of azobenzene, we performed state-averaged complete active space self-consistent field166, 167, 168 calculation with equal weights for and (SA2-CASSCF). All geometry optimization calculations were performed using the cc-pVQZ basis set with JK density fitting169, 170, 171 in the Molpro172, 173, 174 package.
@@
The core ionization energies and XPS spectra were computed using the efficient implementation of MR-ADC in the Prism program.99 As discussed in Section 3, Prism was interfaced with PySCF to obtain the CASSCF orbitals, one- and two-electron integrals, and reference reduced density matrices. All MR-ADC calculations employed the correlation-consistent core-polarized cc-pwCVQZ basis set for carbon and nitrogen175, 176, 177 and the cc-pVQZ basis for hydrogen and iron. Density fitting was used for the reference CASSCF and excited-state MR-ADC methods utilizing the cc-pwCVQZ-JKFIT169, 171 and cc-pwCVQZ-RIFIT169, 178 auxiliary basis sets, respectively. The MR-ADC calculations were performed using the and parameters to remove linearly dependent semiinternal and double excitations, respectively.37 Scalar relativistic effects were incorporated using the spin-free exact-two-component (X2C) Hamiltonian.179, 180, 156, 181, 182 The experimental spectra were digitized using WebPlotDigitizer.183
@@ 5.1 Substituted Ferrocenes
We first apply the efficient MR-ADC implementation developed in this work to three ferrocene derivatives with the ethyl (Et), vinyl (V), or ethynyl (E) substituent (R) in one of the cyclopentadienyl (Cp) rings (Figure 2). The redox potentials of these molecules are highly sensitive to the R group, making them attractive surface modifiers for metals and semiconductors in memory devices, electrochemical sensors, batteries, and fuel cells.185, 186, 187, 188, 189 Meanwhile, the properties of ferrocene and its derivatives are known to be strongly influenced by electron correlation effects and have been the focus of many theoretical studies.190, 191, 192, 193, 194, 195, 196, 197, 198, 199, 200, 201, 202 The carbon K-edge XPS spectra of EtFC, VFC, and EFC in the gas phase have been measured experimentally,184 allowing us to test the performance of MR-ADC for these challenging systems.
Figure 2a shows the experimental C K-edge XPS spectrum of ethyl-ferrocene (EtFC)184 along with the theoretical spectra calculated at the MR-ADC(2) and MR-ADC(2)-X levels of theory. The experimental XPS spectrum shows a broad feature with a maximum at 290.0 eV and a shoulder around 290.5 eV. The MR-ADC(2) and MR-ADC(2)-X calculations reproduce the experimental spectrum quite well when supplied with 0.20 eV broadening. The best agreement with the experiment is shown by MR-ADC(2)-X, which underestimates the band maximum by 0.4 eV and accurately describes the relative energy of the shoulder feature. MR-ADC(2) overestimates the band maximum by 1.4 eV and exhibits a somewhat more pronounced shoulder.
@@
[…]
Finally, we consider the experimental184 and simulated C K-edge XPS spectra of ethynyl-ferrocene (EFC) shown in Figure 2c. Similar to EtFC, the EFC core ionization spectra exhibit a strong peak followed by a shoulder feature 0.5 eV higher in energy. In addition, a weak shoulder is again observed at 289.5 eV, similar to the one in the VFC spectrum. As for EtFC and VFC, MR-ADC(2)-X has the best agreement with the experiment, slightly underestimating the intensity of shoulder at 290.5 eV. In contrast to EtFC where the 290.5 eV shoulder was assigned to the C atom of substituent, the MR-ADC(2)-X calculations suggest that this feature in the EFC spectrum originates from the Cp carbon directly bonded to the ethynyl group. These results suggest that ethynyl acts as an acceptor, depleting the electron density on the Cp carbon adjacent to the substituent and weakening the screening of corresponding core-ionized state. Indeed, the analysis of CASSCF natural orbitals (Figure S4) shows a significant occupation of in-plane and out-of-plane -antibonding ethynyl orbitals, indicating noticeable back-donation from Cp to the ethynyl group. MR-ADC(2) shows a larger error in the peak maximum and predicts a more intense shoulder than what is observed in the experiment as a result of overestimating the relative energy of core-hole states in the substituent. As for VFC, the MR-ADC methods do not predict a weak shoulder at 289.5 eV, which was assigned to the terminal carbon of ethynyl group in the experimental study.184 Although this error may be associated with insufficient description of orbital relaxation effects, it may also originate from the lack of vibrational effects in the simulation and requires further study.
Overall, our results demonstrate that the efficient MR-ADC implementation developed in this work can routinely simulate the K-edge XPS spectra of transition metal complexes with large basis sets. Specifically, the MR-ADC(2) and MR-ADC(2)-X calculations reported here were performed correlating all electrons in 1412 (EFC), 1472 (VFC), and 1532 (EtFC) molecular orbitals. Using a single Intel Xeon Gold 6148 computer node with 40 CPUs, the wall time of MR-ADC(2) and MR-ADC(2)-X simulations did not exceed 4 and 40 hours, respectively, after completing the reference CASSCF calculations. Consistent with our earlier benchmarks,39, 40 the MR-ADC(2)-X results are in very good agreement with the experimental gas-phase XPS spectra.